OpenStudy (anonymous):

P(6, 2) or 6P2???. if you are not going help me get to the answer or go through the steps to get to the answer please dont comment

OpenStudy (amistre64):

theres a formula that defines it, can you type out the formula?

OpenStudy (amistre64):

or are you asking about notation?

OpenStudy (anonymous):

need help Evaluating need to learn the steps on how to do it

OpenStudy (anonymous):

\[P(6,2)={}_6P_2=\frac{6!}{(6-2)!}=\frac{6!}{4!}=\frac{6\times5\times4\times3\times2\times1}{4\times3\times2\times1}=6\times5=\cdots\]